OpenStudy (anonymous):

Write y = -4x2 +16x -14 in vertex form.

OpenStudy (anonymous):

y=-4(x^2-4x+14/4) y=-4(x^2-4x+4-4+14/4) y=-4(x-2)^2+16-14 y=-4(x-2)^2 + 2

OpenStudy (anonymous):

thanks!